ASSIGNING COMPLEX VIBRATION-TUNNELING SPECTRA USING FRANCK-CONDON FINGER-PRINTS

\underline{EDUARDO BERRIOS}, PRAVEEN SUNDARADEVAN and MARTIN GRUEBELE, Department of Chemistry, University of Illinois at Urbana-Champaign, Urbana, Il 61801.

We propose Franck-Condon fingerprinting as a method to assign complex vibration-tunneling spectra. The B̃ electronic excited state of Thiophosgene (SCCl_{2}) is our prototype case. Assignment of its spectrum has proven to be difficult because large tunneling splittings and near-degenerate vibrational frequencies. The dispersed fluorescence spectrum of each unkown vibration-tunneling state reveals its wavefunction character onto the known SCCl_{2} vibrational progressions in the ground states. This Franck-Condon fingerprint allows us to assign several vibration-tunneling states of SCCl_{2} in the B̃ electronic excited state. These assigned transitions could be fitted by an effective vibration-tunneling Hamiltonian within measurement uncertainty.